# Giovanni Compiani Ilya Morozov Stephan Seiler

### v2

@@ 4.2.3 Relevance of Different Data Types
Lastly, we analyze which types of unstructured data yield the largest improvement in fit in each category. It would be misleading to report only the best-performing model in each category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit. We therefore follow the statistical literature on model selection and report Akaike weights for each data type (Burnham and Anderson, 2004). Formally, we define $\Delta_{i}=AIC_{i}-AIC_{min}$ for each estimated specification $i$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category. We then compute the Akaike weight for a data type $d\in\{\text{images},\text{titles},\text{descriptions},\text{reviews}\}$ as follows:
$w_{d}=\frac{\sum_{\{r:d(r)=d\}}\exp(-\Delta_{r}/2)}{\sum_{i}\exp(-\Delta_{i}/2)}$ (2)
@@
The numerator sums over specifications using data type $d$, while the denominator sums over all estimated specifications. We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data. [^25] More precisely, in large samples, $w_{d}$ reflects the probability that this class of models is, in fact, the best model for the data in the sense of Kullback-Leibler information (Burnham and Anderson, 2004, p.272).
@@
Figure 8: Akaike Weights of different types of unstructured data across 40 product categories. These weights reflect the relative importance of data types for predicting substitution patterns in the data.
@@
Importantly, many of these results are difficult to predict ex ante. For example, while we might expect visual product features to be most relevant in clothing categories, where consumers care about visual design, only one of the four clothing categories (“Clothing Underwear”) has an image weight exceeding those of texts. By contrast, estimating demand for “Tops & Blouses” requires product descriptions, for “Shirts” relies on customer reviews, and for “Activewear” it works best with either descriptions or reviews. Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in categories of Nintendo and PC video games, images contain significantly more information about substitution than text.
@@
These results highlight an important lesson: researchers may not be able to predict ex ante which data types will best capture substitution.
5 Discussion
5.1 A Practitioner’s Guide for Using Text and Images
In this section, we summarize our key results and offer guidelines for researchers on how to incorporate unstructured data into demand estimation.
@@ 5.1 A Practitioner’s Guide for Using Text and Images
The main takeaway from our analyses is that text and image data contain information about substitution patterns, meaning these unstructured data can be useful for demand estimation. At the same time, our cross-category analysis in Section 4 highlights that researchers may not be able to predict ex ante which data type will best capture substitution in a given product category. Therefore, we recommend that researchers collect various types of unstructured data and perform model selection as we did in Sections 3 and 4. Specifically, researchers can first identify the lowest-$AIC$ set of principal components for each combination of data type and embedding model, and then determine which of these selected specifications minimizes $AIC$ across all data-model combinations. This selection rule performed well in our experiment, where it identified the model with the best counterfactual second-choice predictions.
In our experimental data in Section 3, we found that a text-based model produces the best counterfactual predictions, and that extending this model with observed attributes or image data did not significantly improve model fit. Nevertheless, in other contexts, researchers may still want to check whether incorporating observed product attributes can capture additional dimensions of product differentiation. This can be done by searching for a set of observed attributes that, when added to the selected model, further reduce $AIC$. A similar search can be conducted over additional text- and image-based principal components to determine whether including them improves model fit.
To conduct inference based on the estimates from our approach, researchers must account for model selection to obtain valid confidence intervals. One potential solution is to perform sample-splitting: researchers can use one randomly selected subset of the data for model selection and the other for estimating the covariance matrix of the estimates. [^26] For a detailed discussion of sample-splitting, see Wasserman and Roeder (2009). Taylor and Tibshirani (2015) provide examples of how researchers can construct corrected p-values and confidence intervals for specific estimation problems.
We mostly abstracted away from the issue of price endogeneity. Endogeneity concerns do not arise in our experiment because we randomize prices, whereas in the analysis of Amazon purchases, we rely on product fixed effects to capture product-level unobservables (e.g., unobserved quality) that may correlate with prices. In practice, researchers are often concerned that prices may be correlated with unobserved demand shocks that vary across markets and over time. To address this, researchers can combine our approach with existing methods for handling price endogeneity based on instrumental variables.
@@ 5.2 Future Research Directions
While we are not the first to use text and image data in demand estimation, our paper proposes and validates a concrete method for leveraging such data to flexibly estimate substitution patterns – a critical input for answering many empirical questions of interest. We do not claim that ours is the only or the best way to use unstructured data for this purpose. In fact, several questions remain open.
First, other demand modeling approaches might be able to better take advantage of text and image data than the random coefficients logit model. In an earlier version of this paper, we tested a pairwise combinatorial logit model from Koppelman and Wen (2000), which generalizes the logit model by allowing each product pair to have its own correlation in utilities. We allowed these correlations to depend on the distance between products in the space of text and image embeddings. This model performed significantly worse than our current approach of incorporating principal components into the standard random coefficients logit model, which highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance. [^27] Researchers can also try relaxing the parametric assumptions of standard demand models. For example, with small number of products, one could incorporate principal components into a nonparametric demand model (Compiani, 2022).
@@
Second, our approach can be easily extended to newer, more advanced ML models that might better extract substitution patterns from texts and images. The approach we propose in this paper is modular in the sense that researchers can easily try alternative embeddings and test whether they improve counterfactual second-choice predictions. To facilitate such tests, we make our experimental dataset and estimation code publicly available. [^28] Replication codes and experimental data are available in our public repository: github.com/ilyamorozov/DeepLogitReplication. Separately, the Python package for implementing our method is available on PyPI and at: github.com/deep-logit-demand/deeplogit.
Third, researchers could improve upon our approach by constructing “optimal” embeddings specifically designed to improve the counterfactual predictions of interest. [^29] For example, Lee (2024) shows that fine-tuning LLMs helps researchers to better capture brand intercepts. Our current method relies on text and image models that were trained on datasets unrelated to our specific application, meaning they may not capture the most relevant dimensions of product differentiation. This suggests potential for improvement: researchers could fine-tune existing deep learning models or train custom ones to optimize for a specific objective, such as improving counterfactual predictions. It would be interesting to assess how much these tailored demand models improve upon pre-trained models and whether they can do so without significantly increasing the computational burden.
Lastly, our approach could extend to other forms of unstructured data, such as music or videos, potentially opening doors to studying product differentiation and competition in markets where this would otherwise be challenging. One could also leverage tools from the explainable AI literature to construct interpretable embeddings, which would help tackle questions such as optimal product design. [^30] See Sisodia et al. (2024) for a detailed discussion of related ideas.
@@ 6 Conclusion
In this paper, we demonstrated how researchers can incorporate unstructured text and image data in demand estimation to better recover substitution patterns. Our proposed estimation approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard discrete choice model. Using experimental data, we showed that our model outperforms standard attribute-based models in counterfactual predictions of second choices. We further validated our method with e-commerce data across dozens of categories and found that text and image data consistently help identify close substitutes within each category.
@@ Appendix A Text Processing Steps
Before applying our text models, listed in Section 2.1.2, we pre-process our text data as follows. Working with product titles is straightforward because each title is a short text that typically includes only 10-15 words. When assessing similarity based on product descriptions, we merge the text from all bullet points, and apply our models to the merged text. For customer reviews, we transform the text of each review into a separate vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
@@
Appendix B Choice Survey: Sanity Checks
@@
Appendix C Changes in Text and Images Over Time
As discussed in Section 2.3, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product. To validate this assumption, we examine whether text and images change over time.
@@ Appendix C Changes in Text and Images Over Time
Images: InceptionV3 -20475.8 40985.6 0.080 -0.009 -9.8%
Images: ResNet50 -20479.9 40991.8 0.080 -0.009 -9.8%
Images: VGG19 -20481.2 40994.4 0.083 -0.006 -6.8%
Images: Xception -20479.1 40990.3 0.086 -0.003 -3.1%
Product Titles: Bag-of-Words Count -20479.3 40992.7 0.085 -0.004 -4.6%
Product Titles: Bag-of-Words TF-IDF -20479.3 40992.7 0.085 -0.004 -4.6%
Product Titles: Sentence Encoder (USE) -20479.1 40992.2 0.080 -0.009 -10.5%
Product Titles: Sentence Transformer (ST) -20477.2 40986.5 0.082 -0.006 -7.1%
Descriptions: Bag-of-Words Count -20480.1 40992.2 0.089 -$0$0.2%
Descriptions: Bag-of-Words TF-IDF -20480.2 40992.3 0.089 -$0$0.2%
Descriptions: Sentence Encoder (USE) -20476.2 40984.4 0.075 -0.014 -15.5%
Descriptions: Sentence Transformer (ST) -20475.8 40983.6 0.073 -0.016 -18.0%
Reviews: Bag-of-Words Count -20481.0 40995.9 0.084 -0.005 -5.2%
Reviews: Bag-of-Words TF-IDF -20480.9 40995.7 0.084 -0.005 -5.1%
Reviews: Sentence Encoder (USE) -20472.0 40975.9 0.069 -0.020 -22.3%
Reviews: Sentence Transformer (ST) -20473.3 40978.6 0.072 -0.017 -18.8%
Panel C. Attribute-Based Mixed Logit Models
Price -20485.9 41003.9 0.088 -0.001 -0.8%
Pages -20484.9 41001.7 0.083 -0.006 -6.8%
Year -20484.7 41001.4 0.087 -0.002 -2.0%
Genre -20483.3 41000.7 0.081 -0.008 -8.9%
Price & Pages -20478.9 40991.9 0.077 -0.012 -13.0%
Price & Year -20484.1 41002.2 0.088 -0.001 -1.4%
Price & Genre -20483.2 41002.5 0.080 -0.008 -9.4%
Pages & Year -20478.3 40990.7 0.077 -0.011 -12.9%
Pages & Genre -20481.6 40999.2 0.076 -0.013 -14.9%
Year & Genre -20483.3 41000.7 0.081 -0.008 -8.9%
Price, Pages, & Year -20478.3 40990.7 0.077 -0.011 -12.9%
Price, Pages, & Genre -20478.9 40991.9 0.077 -0.012 -13.0%
Price, Year, & Genre -20482.2 41002.4 0.079 -0.010 -11.2%
Pages, Year, & Genre -20478.3 40990.7 0.077 -0.011 -12.9%
Price, Pages, Year, & Genre (All Attr.) -20476.9 40993.9 0.076 -0.013 -14.4%
@@
Table A1: Model validation results. The table shows in-sample fit on first-choice data and counterfactual performance on second-choice data for all specifications considered in Figures 3 and 4 (see Section 3.2 for detailed description of these models).

### v3

@@ 4.3 Relevance of Different Data Types
Lastly, we analyze which types of unstructured data yield the largest fit improvements. It would be misleading to report only the best-performing model in each category, as multiple text or image models may achieve similar $AIC$ improvements over plain logit. We therefore follow the statistical literature on model selection and report Akaike weights (Burnham and Anderson, 2004). Formally, for each data type $d$ (images, titles, descriptions, or reviews) we compute the Akaike weight $w_{d}$ as
$w_{d}=\frac{\sumop\displaylimits_{\{r:d(r)=d\}}\exp(-{}_{r}/2)}{\sumop\displaylimits_{i}\exp(-{}_{i}/2)}$ (3)
@@
where $i$ indexes estimated specifications, and $_{i}=AIC_{i}-AIC_{min}$ where $AIC_{min}$ is the lowest $AIC$ across all estimated specifications in that product category. The numerator in (3) sums over estimated specifications using data type $d$, while the denominator sums over all specifications estimated during model selection. We interpret $w_{d}$ heuristically as the posterior probability that the mixed logit model based on data type $d$ is the best model given the data. [^26] More precisely, in large samples, $w_{d}$ reflects the probability that this class of models is, in fact, the best model for the data in the sense of Kullback-Leibler information (Burnham and Anderson, 2004, p.272).
@@
Figure 8: Akaike Weights of different types of unstructured data across 40 product categories. These weights reflect the relative importance of data types for predicting substitution patterns in the data. We compute Akaike weights using the formula in (3).
@@
Importantly, many of these results are difficult to predict in advance. For example, while we might expect visual features to be most relevant in clothing categories, where consumers care about visual design, only two of the five clothing categories (“Activewear” and “Sleep”) assign more weight to images than to text. In contrast, for “Tops & Blouses” and “Underwear,” product titles are most predictive of substitution, whereas for “Shirts,” reviews perform best. Similarly, while we may expect descriptions and reviews of video games to be as informative as those for books, the data strongly suggests that in the category of video games for Xbox, images contain substantially more information about substitution patterns than text.
@@
These results highlight that researchers cannot reliably predict in advance which data types will best capture substitution. Therefore, in practice, it is important to collect different data types and use model selection to identify the one that is most predictive of substitution.
5 A Practitioner’s Guide for Using Text and Images
In this section, we provide guidance for practitioners interested in applying our method. We highlight several practical lessons from Sections 3 and 4, clarify which counterfactuals our approach can and cannot accommodate, and suggest promising directions for future research.
5.1 Data Choice and Model Selection
@@
The main takeaway from our results is that text and image data contain valuable information about substitution patterns, making these unstructured data useful for demand estimation. At the same time, our multi-category analysis in Section 4 shows that researchers may not be able to predict in advance which data type will best capture substitution in a given category. We therefore recommend collecting various types of unstructured data and performing model selection as in Sections 3 and 4. Algorithm 1 in Section 2 provides a practical heuristic for model selection: in our experiment, it successfully identified the specification that delivered the best counterfactual second-choice predictions.
A natural question is whether researchers should add observed attributes to our approach when they are available. In our application, adding observed attributes to the selected review-based model did not improve fit (see Section 3.3). One explanation is that unstructured data may already encode the choice-relevant attributes, rendering structured attributes redundant. Nevertheless, in other applications, researchers may still want to test whether including observed attributes helps estimate substitution patterns. This can be done by checking whether adding some subset of observed attributes to the selected model with unstructured data further reduces $AIC$. When many attributes are available, researchers can reduce their dimensionality using PCA, as in Section 4.2.
5.2 Counterfactual Analysis and its Boundaries
We interpret embeddings as capturing time-invariant product aspects (e.g., attributes, functional benefits, and visual designs). This assumption is reasonable given that in our application, product images, titles, and descriptions rarely change over time (see Appendix F). Thus, our method is well-suited for a wide range of counterfactuals that require estimating consumer responses to price changes conditional on the other product attributes, including optimal pricing, simulating merger effects on prices, evaluating corrective taxes, or estimating markups (Berry and Haile, 2024). In such contexts, researchers can evaluate counterfactual market shares and prices just as in standard demand models, assuming that embeddings—and the resulting demand functions—do not change. Moreover, as usual, researchers can quantify the welfare effects of price changes by calculating the area under the demand curve (Small and Rosen, 1981). The fact that principal components have no direct interpretation does not preclude such analysis.
Our approach can also be applied to study consumer responses to changes in product availability, for example, to quantify consumer surplus from new products or study how firms optimize assortments. In retrospective analyses, where the introduction of new products is observed in the data as in Petrin (2002), researchers can use our approach “as is” knowing that it performs well at predicting consumer responses to product removals. By contrast, evaluating a hypothetical new product that is not yet in the data poses an additional challenge: researchers need to obtain images and textual descriptions of this product from pre-launch information.
In some applications, embeddings may change in counterfactual scenarios. For example, if customer reviews mention value relative to price (e.g., “Great quality for the price!” or “Good product, but overpriced.”), then the reviews and the resulting text embeddings will shift when prices change in counterfactuals. Researchers can diagnose this issue by checking how common such reviews are in their data and, if necessary, model how price changes affect embeddings.
Another example is merger simulations: if firms redesign or reposition products after merging, this may alter embeddings and hence substitution patterns. If this is a first-order concern, researchers may want to specify a supply-side model in which firms choose both prices and non-price characteristics (Fan, 2013).
That said, our method does allow for correlation between embeddings and variables that change in counterfactuals. For instance, embeddings may be correlated with prices if they capture unobserved product quality. This does not rule out counterfactuals where prices change (e.g., due to tariffs), since such changes do not alter the product’s intrinsic quality.
@@ 5.3 Future Research Directions
Our paper shows that text and image data contain valuable and easily extractable information for estimating substitution patterns. This finding opens several promising directions for future research.
@@
First, newer ML models might extract substitution patterns from texts and images more effectively. For example, Qwen3, OpenAI, and Gemini text embeddings perform robustly well across diverse natural language tasks and may thus generalize to demand estimation (Lee et al., 2025; Zhang et al., 2025). Because our method is modular, researchers can easily try alternative embeddings and test whether that improves performance. In addition, researchers can fine-tune existing models to construct embeddings optimized for counterfactual predictions. An open question is how much fine-tuning improves upon pre-trained models and whether these gains come without excessive computational costs. To facilitate future research on alternative embedding models, we make our experimental dataset and estimation code publicly available. [^27] Replication codes and experimental data are available in our public repository: github.com/ilyamorozov/DeepLogitReplication. Separately, the Python package for implementing our method is available on PyPI and at: github.com/deep-logit-demand/deeplogit.
Second, future research should provide a formal treatment of how to perform inference in a way that accounts for the uncertainty in the model selection step. One possibility would be to use sample-splitting as in Wasserman and Roeder (2009) and Taylor and Tibshirani (2015), randomly selecting a subset of the data for model selection and using the other for estimation and inference based on the selected model.
Third, other demand models may leverage text and image data more effectively than the random coefficients logit model. In an earlier version of this paper, we tested the pairwise combinatorial logit model of Koppelman and Wen (2000), which allows each product pair to have its own utility correlation based on distance in the embedding space. The model, however, performed significantly worse than our mixed logit, not only when using unstructured data but also with observed attributes, where we let utility correlations depend on the distance between products in the attribute space. This highlights how the functional form assumptions of the demand model can significantly impact counterfactual performance and suggests that combining unstructured data with more flexible demand models, such as that in Compiani (2022), may be a fruitful direction for future research. We note that this point applies more generally to any discrete choice model and is not specific to our method.
Finally, because our focus is on recovering substitution patterns, we largely abstract away from price endogeneity. Endogeneity is not a concern in our experiment because we randomize prices. In the Amazon application, we rely on product fixed effects to capture unobserved attributes (e.g., quality) that may correlate with prices. More generally, researchers may worry that prices correlate with unobserved demand shocks that vary across markets and over time. To address this, one could combine our approach with existing methods for handling price endogeneity using instrumental variables (Goolsbee and Petrin, 2004; Berry et al., 2004). [^28] For example, one could estimate product-market fixed effects and random coefficient variances via maximum likelihood, and then estimate a two-stage least squares regression of the estimated fixed effects on prices and principal components. The principal components of competing products could serve as excluded instruments for prices.
@@ 6 Conclusion
In this paper, we demonstrate how researchers can incorporate unstructured text and image data in demand estimation to recover substitution patterns. Our approach extracts low-dimensional features from product images and textual descriptions, integrating them into a standard mixed logit model with random coefficients. Using experimental data, we show that our approach outperforms standard attribute-based models in counterfactual predictions of second choices. We further validate our method with e-commerce data across dozens of categories and find that text and image data consistently help identify close substitutes within each category.
@@ Appendix A Text Processing Steps
Before applying our text models, listed in Section 2, we pre-process our text data as follows. Working with product titles is straightforward because each title is a short text that typically includes only 10-15 words. When assessing similarity based on product descriptions, we merge the text from all bullet points, and apply our models to the merged text. For customer reviews, we transform the text of each review into a separate vector of word occurrences or an embedding, and we average these vectors or embeddings across all reviews of a given product.
@@
Appendix B Book Selection Procedure
Which books we show to participants can significantly affect the performance of different demand models. Our general strategy was to select a set of books that does not bias our model comparisons in favor of any particular specification. To this end, we aimed to maximize variation in choice-relevant structured attributes as well as in text and image embeddings.
We chose books from three genres—Mystery, Fantasy, and Self-Help—and identified 20 books of each genre from Amazon’s bestseller lists. We included books from three different genres in order to generate meaningful variation in this structured attribute, which we anticipated would be predictive of substitution patterns.
To avoid arbitrary book selection within each genre, we implemented the following algorithm. We considered all possible combinations of ten books such that the three genres are roughly equally represented. Among these, we choose the combination of books that maximized the variance of text and image embeddings. [^29] Operationally, we computed the variance of image embeddings (using the VGG19 model) and the variance of text embeddings (using the USE model), then averaged the two. We performed a brute-force search over all possible sets of ten books and selected the set that maximized this average variance. The final set, shown in Figure 3, includes four mystery books, three fantasy books, and three self-help books.
Appendix C Choice Survey: Sanity Checks
@@
Appendix D Computing Second-Choice Diversion Ratios
To compute $RMSE$ in (2), we need to calculate predicted second-choice diversion ratios $\hat{s}_{j\rightarrow k}$ and their data counterparts $s_{j\rightarrow k}$. Recall that both $s_{j\rightarrow k}$ and $\hat{s}_{j\rightarrow k}$ reflect the probability that the consumer chooses book k in the second choice task conditional on having chosen book j in the first choice task. Using the analogy principle, we compute empirical diversions $s_{j\rightarrow k}$ using a simple frequency estimator:
$s_{j\rightarrow k}=\frac{\sumop\displaylimits_{i=1}^{N}\mathbf{1}\{y_{i}^{(2)}=k,y_{i}^{(1)}=j\}}{\sumop\displaylimits_{i=1}^{N}\mathbf{1}\{y_{i}^{(1)}=j\}}\quad\text{for}\quad j\neq k,$ (4)
where $y_{i}^{(1)}$ and $y_{i}^{(2)}$ are participant $i$’s first and second choices. By contrast, to compute diversions $\hat{s}_{j\rightarrow k}$ predicted by a given demand model, we use the identity
$Pr(y_{i}^{(2)}=k|y_{i}^{(1)}=j;w_{i})=\frac{s_{k}^{(j)}(w_{i})-s_{k}(w_{i})}{s_{j}(w_{i})},$ (5)
where $s_{j}(w_{i})$ and $s_{k}(w_{i})$ are the unconditional choice probabilities of books $j$ and $k$, $s_{k}^{(j)}(w_{i})$ is the probability of choosing book $k$ after book $j$ is removed from the choice set, and $w_{i}$ is a vector of prices $(\text{price}_{i1},\dots,\text{price}_{iJ})$ and rankings $(\text{rank}_{i1},\dots,\text{rank}_{iJ})$ that participant $i$ encounters in the experiment. We average diversions in (5) across all participants to compute $\hat{s}_{j\rightarrow k}$:
$\hat{s}_{j\rightarrow k}=\frac{1}{N}\sumop\displaylimits_{i=1}^{N}\frac{s_{k}^{(j)}(w_{i})-s_{k}(w_{i})}{s_{j}(w_{i})}.$ (6)
Appendix E Attributes Collected for Electronics Products
For each of the four electronics categories described in Section 4.1, we collect standard attributes from Amazon product detail pages. Specifically, we include all specifications listed in the product descriptions and, when available, those in the “Technical Details”’ section. Below is the complete list of all attributes for each category:
  1. 1.
    Tablets: brand, model, memory, RAM, processor speed, number of cores, screen size, maximum resolution, charging time, battery life, front camera, back camera, front camera megapixels, back camera megapixels, warranty, whether the product comes with a case, and whether it includes an Amazon Kids subscription.
  2. 2.
    Monitors: brand, screen size, maximum resolution, refresh rate, blue light filter, frameless design, tilt adjustment, height adjustment, flicker-free display, built-in speaker, wall-mountable option, curved screen, and adaptive sync.
  3. 3.
    Memory Cards: brand, pack size, micro card, flash memory type, capacity, read speed, write speed, speed class, UHS speed class, device compatibility (smartphone, computer, camera, laptop, tablet), X-ray proof, temperature proof, waterproof, shockproof, magnetic proof, and whether an adapter is included.
  4. 4.
    Headphones: brand, color, connectivity, number of eartip sets, battery life on a single charge, battery life with charging case, deep bass feature, tangle-free design, waterproof, sweat-proof, IPX rating, whether the product comes with a case, a microphone, and a noise reduction option.
Appendix F Changes in Text and Images Over Time
As discussed in Section 2, we treat text and image embeddings, as well as the principal components extracted from them, as being fixed over time for a given product. To validate this assumption, we examine whether text and images change over time.
@@ Appendix F Changes in Text and Images Over Time
Figure A4: Share of variance of embeddings explained by principal components.
[…]
@@
Images: InceptionV3 -20475.2 40990.5 0.085 -0.006 -7.0%
Images: ResNet50 -20479.9 40997.8 0.083 -0.009 -9.4%
Images: VGG19 -20481.2 41000.4 0.087 -0.005 -5.0%
Images: Xception -20479.1 40996.3 0.089 -0.002 -2.1%
Product Titles: Bag-of-Words Count -20478.7 40995.5 0.084 -0.007 -7.9%
Product Titles: Bag-of-Words TF-IDF -20478.7 40995.5 0.084 -0.007 -7.9%
Product Titles: Sentence Encoder (USE) -20481.2 40998.3 0.087 -0.004 -4.3%
Product Titles: Sentence Transformer (ST) -20477.1 40992.3 0.085 -0.007 -7.3%
Descriptions: Bag-of-Words Count -20475.5 40988.9 0.085 -0.006 -6.9%
Descriptions: Bag-of-Words TF-IDF -20475.8 40989.6 0.085 -0.006 -6.5%
Descriptions: Sentence Encoder (USE) -20474.2 40986.4 0.076 -0.015 -17.0%
Descriptions: Sentence Transformer (ST) -20474.9 40987.8 0.075 -0.016 -18.1%
Reviews: Bag-of-Words Count -20479.6 40997.1 0.082 -0.009 -9.8%
Reviews: Bag-of-Words TF-IDF -20479.4 40996.9 0.082 -0.009 -10.1%
Reviews: Sentence Encoder (USE) -20472.0 40981.9 0.070 -0.021 -23.0%
Reviews: Sentence Transformer (ST) -20473.3 40984.6 0.073 -0.019 -20.4%
Panel C. Mixed Logit with Observed Attributes
Price -20485.9 41003.9 0.091 -0.000 -0.0%
Pages -20484.9 41001.7 0.086 -0.005 -5.5%
Year -20484.7 41001.4 0.091 0.000 0.1%
Genre -20483.3 41000.7 0.084 -0.007 -7.7%
Price & Pages -20478.9 40991.9 0.081 -0.010 -11.4%
Price & Year -20484.1 41002.2 0.092 0.000 0.3%
Price & Genre -20483.2 41002.5 0.086 -0.005 -5.2%
Pages & Year -20478.3 40990.7 0.081 -0.011 -11.7%
Pages & Genre -20481.6 40999.2 0.081 -0.010 -11.4%
Year & Genre -20483.3 41002.7 0.084 -0.007 -7.7%
Price, Pages, & Year -20478.3 40992.7 0.081 -0.011 -11.7%
Price, Pages, & Genre -20478.9 40995.9 0.081 -0.010 -11.4%
Price, Year, & Genre -20482.2 41002.4 0.084 -0.007 -7.4%
Pages, Year, & Genre -20478.3 40994.7 0.081 -0.011 -11.7%
Price, Pages, Year, & Genre (All Attr.) -20476.9 40993.9 0.078 -0.013 -14.1%
@@
Table A1: Model validation results. The table shows in-sample fit on first-choice data and counterfactual performance on second-choice data for all specifications considered in Figure 2 (see Section 3.2 for detailed description of these models).
Book Plain Logit Mixed Logit with Attributes Mixed Logit with Texts
Don’t Believe 3.7% 6.7% 8.9%
Art of Letting Go 3.0% 4.3% 5.6%
The Ritual 3.1% 1.6% 2.5%
The Inmate 5.0% 4.4% 4.3%
Ashes & Star 2.7% 1.3% 2.0%
Court of Ravens 3.3% 3.5% 2.5%
Serpent & Wings 3.4% 2.3% 2.3%
Please Tell Me 6.9% 6.1% 5.6%
The Housemaid 5.4% 5.2% 4.0%
Table A2: Merger Simulation Results (Dopamine Detox). The table shows predicted relative price increases resulting from a simulated merger between Dopamine Detox and each other book respectively. For each simulated merger between Dopamine Detox and another book (first column), the table reports the expected average price increase for the two merging books across three estimated demand models.
Category Data Type Model Type AIC Diversion to Closest Substitute
1. Clothing Active Images VGG16 -25.1 11.5%
2. Clothing Shirts Reviews USE -11.9 12.5%
3. Clothing Underwear Titles TFIDF -16.3 45.5%
4. Clothing Sleep Images Inceptionv3 -13.1 32.5%
5. Clothing Tops & Blouses Titles USE -12.3 30.4%
6. Electronics Cables Reviews TFIDF -16.7 20.1%
7. Electronics Accessories Descriptions COUNT -19.6 19.0%
8. Electronics Keyboards Descriptions ST -19.5 20.0%
9. Electronics Memory Cards Images VGG16 -96.8 33.4%
10. Electronics Tablets Descriptions ST -111.5 18.2%
11. Electronics Monitors Images Resnet50 -22.4 19.9%
12. Electronics Headphones Images VGG19 -10.7 6.1%
13. Electronics Media Players Images VGG19 -90.2 23.4%
14. Groceries Water Titles USE -19.1 19.0%
15. Groceries Coffee Descriptions TFIDF -9.5 31.8%
16. Groceries Tea Images Xception -17.1 37.4%
17. Groceries Chips Descriptions ST -5.6 19.2%
18. Household Aromatherapy Descriptions COUNT -11.2 17.6%
19. Household Batteries Images VGG16 -10.7 -0.6%
20. Household Trash Bags Descriptions TFIDF -16.2 40.1%
21. Household Paper Towels Titles TFIDF -16.9 33.1%
22. Home Sheets & Pillowcases Images Resnet50 -21.3 6.8%
23. Bedroom Beds Descriptions USE -11.2 26.0%
24. Bedroom Mattresses Reviews ST -12.9 23.5%
25. Kitchen Food Storage Images VGG19 -20.5 28.9%
26. Office Folders Descriptions ST -17.4 31.9%
27. Office Paper Reviews ST -13.5 21.0%
28. Office Markers Images Inceptionv3 -13.6 19.8%
29. Office Pens Images VGG19 -13.5 31.6%
30. Office Printer Supplies Titles USE -15.1 34.1%
31. Pet Cat Food Descriptions COUNT -11.1 36.2%
32. Pet Cat Litter Titles ST -12.1 34.1%
33. Pet Cat Snacks Reviews COUNT -16.9 29.1%
34. Pet Dog Food Images Inceptionv3 -11.7 24.5%
35. Pet Dog Treats Titles USE -18.2 38.0%
36. Game Consoles Nintendo Images Resnet50 -6.8 11.9%
37. Video Games Nintendo Descriptions ST -23.5 13.3%
38. Video Games PC Titles COUNT -9.5 20.7%
39. Video Games PS4 Descriptions ST -34.7 32.2%
40. Video Games Xbox Images Xception -74.7 28.9%
Averaged -23.3 24.6%
Table A3: Estimation results across 40 categories in Comscore data. For each category, the table shows the selected model and data type yielding the lowest AICAIC and the $AIC$ improvement relative to plain logit. The last column shows the increase in the predicted diversion ratios to the closest substitutes, $\max_{k}\hat{s}_{j\rightarrow k}$, averaged across products $j$, relative to plain logit (in percentage points).
[…]
Category: Tablets
Mixed Logit with Attributes 4293.2 -54.9
Mixed Logit with Unstructured Data 4236.6 -111.5
Category: Monitors
Mixed Logit with Attributes 1376.4 -12.3
Mixed Logit with Unstructured Data 1366.4 -22.4
Category: Memory Cards
Mixed Logit with Attributes 2709.3 -72.6
Mixed Logit with Unstructured Data 2685.1 -96.8
Category: Headphones
Mixed Logit with Attributes 9882.0 -4.5
Mixed Logit with Unstructured Data 9875.8 -10.7
Table A4: AICAIC Comparison of Our Approach and Mixed Logit with Attributes
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.330 0.381 0.301 0.312 0.283 0.279 0.286
Fire HD 8 0.225 0.000 0.229 0.175 0.178 0.165 0.162 0.166
Fire HD 10 0.386 0.340 0.000 0.314 0.320 0.297 0.291 0.295
Fire 7 Kids 0.123 0.105 0.128 0.000 0.100 0.091 0.089 0.091
iPad 10.2 0.153 0.129 0.148 0.122 0.000 0.113 0.112 0.114
Fire HD 8 Kids 0.043 0.038 0.049 0.035 0.035 0.000 0.032 0.033
Dragon Touch 0.022 0.018 0.021 0.017 0.017 0.016 0.000 0.016
iPad 9.7 0.048 0.041 0.044 0.037 0.037 0.035 0.034 0.000
Table A5: Estimated Diversion Ratios for Tablets (Plain Logit). Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.
Fire 7 Fire HD Fire HD Fire 7 iPad Fire HD Dragon iPad
8 10 Kids 10.2 8 Kids Touch 9.7
Fire 7 0.000 0.590 0.162 0.362 0.143 0.033 0.024 0.080
Fire HD 8 0.753 0.000 0.429 0.336 0.200 0.193 0.108 0.116
Fire HD 10 0.058 0.144 0.000 0.079 0.142 0.297 0.022 0.077
Fire 7 Kids 0.127 0.131 0.080 0.000 0.012 0.440 0.293 0.007
iPad 10.2 0.049 0.079 0.141 0.010 0.000 0.015 0.310 0.715
Fire HD 8 Kids 0.003 0.034 0.151 0.200 0.004 0.000 0.188 0.001
Dragon Touch 0.000 0.001 0.000 0.011 0.013 0.019 0.000 0.003
iPad 9.7 0.011 0.020 0.036 0.002 0.485 0.003 0.055 0.000
Table A6: Estimated Diversion Ratios for Tablets (Selected Model: Descriptions ST). Each cell reports the probability of choosing the row product $j$ when the first choice, column product $k$, is removed from the choice set.